OpenStudy (zuha):

find a function rule x -2 , -1 , 0 , 1 , 2 y 9 , 4 , -1 , -6 , -11 please help!!

OpenStudy (anonymous):

first make a graph and plot any three of the points

OpenStudy (zuha):

ok

OpenStudy (zuha):

then?

OpenStudy (anonymous):

we see if it is a straight line, is it?

OpenStudy (zuha):

ok hold on

OpenStudy (zuha):

yes

OpenStudy (anonymous):

awesome, so we know its a line. that means we can make a function starting from the standard function of a line: y = mx+b do you remember how to find the slope between two points? if so, pick any two points and find the slope, if not, just tell me?

OpenStudy (zuha):

-2 , 2

OpenStudy (zuha):

then

OpenStudy (zuha):

hello

OpenStudy (anonymous):

to find the slope I do, the change in y of two points divide by the change in x of two points: i'll pick points, (-2, 9) (-1, 4) looking at the y values I have 9 and 4, the change to go from 9 to 4 is -5 looking at the x values I have -2 and -1, the change to go from -2 to -1 is 1 so my slope is -5/1 or -5 now we know m = -5 lets plug in any point given to us, and the value of m in to the equation y = mx+b and solve for b

OpenStudy (anonymous):

you sent me a message, am I not helping you the way you would like?

OpenStudy (zuha):

no I thought you went away ! sorry!

OpenStudy (anonymous):

^_^

OpenStudy (zuha):

ok, let me check

OpenStudy (zuha):

How do I solve b?

OpenStudy (anonymous):

what point did you pick from what they gave us in the problem and i'll show you?

OpenStudy (zuha):

(0, -1) and (1, -6)

OpenStudy (anonymous):

we only need 1 points, so i'll use (0, -1) y = mx+ b m = -5 x = 0 y = -1 -1 = -5*0 + b solve for b

OpenStudy (zuha):

ok

OpenStudy (zuha):

is b 0?

OpenStudy (anonymous):

-1 = -5*0 + b -1 = 0 + b -1 = b

OpenStudy (zuha):

ohh, so b is -1?

OpenStudy (anonymous):

yep! so we are basically done after this next step. the problem wanted us to find the function of the points. So, we found what m equals and what b equals. just plug in m and b into the equation y = mx+b and the equation becomes the answer

OpenStudy (zuha):

oh, thanks!!

OpenStudy (anonymous):

^_^